OpenStudy (anonymous):

sin(theta/2)-1=0 Find the solutions in the interval [0,2pi]

OpenStudy (xapproachesinfinity):

\(\rm sin(\theta/2)-1=0\Longrightarrow sin(\theta/2)=1\) know you ask what angles give 1 in the range [0. 2pi]

OpenStudy (xapproachesinfinity):

you have pi/2 only correct?

OpenStudy (xapproachesinfinity):

So?

OpenStudy (xapproachesinfinity):

Don't wait there for a direct answer! you need to show some work if you want help

OpenStudy (xapproachesinfinity):

you need to say something did you understand what im doing or not if not say so, so i explain more

OpenStudy (xapproachesinfinity):

Staying quite there is not gonna help you know

OpenStudy (anonymous):

chill, was afk

OpenStudy (anonymous):

only solution is pi.

OpenStudy (xapproachesinfinity):

hhehe i'm chilled! it is actually cold outside so real chill just joking dude! and yes pi is the solution